# Unbiased Region-Language Alignment for Open-Vocabulary Dense Prediction

### v2

@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [57, 13, 38, 23] and image segmentation [11, 52, 29, 53, 4, 28], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [63, 21, 53] for dense prediction have achieved great progress.
Popular VLMs, such as CLIP [42] and EVA-CLIP [45], have exhibited remarkable zero-shot recognition abilities for global image understanding. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [60, 63]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [60, 27, 59, 32] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [48] and MaskEmbed [5], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment. Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [45, 60, 48]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [56]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [64, 48].
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [57, 13, 38, 23] and image segmentation [11, 52, 29, 53, 28]. The success of pre-trained vision-language models like CLIP [42], has further inspired advancements in this area. In open-vocabulary detection, recent studies [13, 47] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [23, 49] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [7, 6, 52, 34] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [54] or the adoption of single-stage frameworks [55]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [4, 51] or mask-level [19] alignment, they are constrained by the scarcity of densely labeled data.
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [17, 37, 42]. By using image-text pairs [44] as training data, methods such as CLIP [42] and ALIGN [17] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [46, 24, 25, 50] attempt to modify the output layers of CLIP. Recent methods [40, 43, 41] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
@@ 3.2 DenseVLM framework
To mitigate this issue, we propose a decoupled alignment strategy that separates the alignment process for foreground and background regions. Specifically, we denote the region features of the training U-VLM as , using the same partitioning grid as the P-VLM. The text embeddings are represented as . By leveraging the region-category relations retrieved by the P-VLM, we establish a one-to-one mapping for the U-VLM to associate the region features and their corresponding category embeddings. To distinguish the semantic regions of foreground and background, we decouple these region-category relations into two distinct groups, following the predefined two category sets: foreground objects Thing () and background regions Stuff (). By selectively contrasting against text embeddings related to irrelevant categories, we guide the model to focus more on the relevant background regions, reducing the impact of irrelevant foreground categories. This selective contrast helps the model capture the distinctive characteristics of background regions, leading to a more accurate separation between foreground and background. The alignment process can be effectively achieved by maximizing the cosine similarity for the region features and the text embeddings. According to Eqn. (2), when , the probability for a specific region is calculated as:
@@ 4.3 Ablation study
Framework design exploration. We conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [48] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction, while introducing foreground bias issues. In contrast, DenseVLM achieves superior performance with a more efficient feature-cropping strategy and decoupled alignment framework, reducing both training time and GPU memory usage.
@@ 4.4 Zero-shot comparisons at different data scales
4.5 Extending to different backbones
4.6 Application to open-vocabulary dense tasks
@@ 5 Conclusions
In this paper, we present DenseVLM, a framework designed to mitigate foreground bias in region-level vision-language alignment. DenseVLM can be seamlessly applied to open-vocabulary object detection and image segmentation tasks, demonstrating consistently clear performance improvements over the baseline models. Furthermore, we validate DenseVLM’s promising scaling ability by exploring efficient region retrieval and decoupled alignment, successfully implementing DenseVLM to the training data from the SA-1B dataset. Overall, DenseVLM offers a generalizable solution for improving dense representations in vision-language models across various backbones, which is essential for open-vocabulary dense prediction tasks.

### v3

# DenseVLM: A Retrieval and Decoupled Alignment Framework for Open-Vocabulary Dense Prediction

@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [68, 15, 48, 29, 22] and image segmentation [13, 63, 38, 64, 5, 35], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [74, 27, 64] for dense prediction have achieved great progress.
Popular VLMs, such as CLIP [52] and EVA-CLIP [55], have exhibited remarkable zero-shot recognition abilities for global image understanding [19, 20, 37, 36, 12]. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [71, 74]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [71, 34, 70, 41] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [59] and MaskEmbed [6], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment, as shown in Fig. 2(d). Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [55, 71, 59]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [67]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [75, 59].
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [68, 15, 48, 29] and image segmentation [13, 63, 38, 64, 35, 33]. The success of pre-trained vision-language models like CLIP [52], has further inspired advancements in this area. In open-vocabulary detection, recent studies [15, 58] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [29, 60] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [8, 7, 63, 43] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [65] or the adoption of single-stage frameworks [66]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, which is essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [5, 62] or mask-level [24] alignment, they are constrained by the scarcity of densely labeled data.
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [21, 46, 52, 26, 47, 2, 57]. By using image-text pairs [54] as training data, methods such as CLIP [52] and ALIGN [21] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [56, 30, 31, 61] attempt to modify the output layers of CLIP. Recent methods [50, 53, 51] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
@@ 3.2 DenseVLM framework
To mitigate this issue, we propose a decoupled alignment strategy that separates the alignment process for foreground and background regions. Specifically, we denote the region features of the training U-VLM as , using the same partitioning grid as the P-VLM. The text embeddings are represented as . By leveraging the region-category relations retrieved by the P-VLM, we establish a one-to-one mapping for the U-VLM to associate the region features and their corresponding category embeddings. To distinguish the semantic regions of foreground and background, we decouple these region-category relations into two distinct groups, following the predefined category sets: foreground objects Thing () and background regions Stuff (). By selectively contrasting against text embeddings related to irrelevant categories, we guide the model to focus more on the relevant background regions, reducing the impact of irrelevant foreground categories. This selective contrast helps the model capture the distinctive characteristics of background regions, leading to a more accurate separation between foreground and background. The alignment process can be effectively achieved by maximizing the cosine similarity for the region features and the text embeddings. According to Eqn. (2), when , the probability for a specific region is calculated as:
@@ 4.3 Ablation study
Framework design exploration. As shown in Tab. 2, we conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [59] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction while introducing foreground bias issues. In contrast, DenseVLM leverages a more efficient feature-cropping strategy and a decoupled alignment framework to achieve superior performance while substantially reducing both training time and GPU memory usage.
@@ 4.4 Zero-shot comparisons at different data scales
4.5 Extending to different backbones
4.6 Application to open-vocabulary dense tasks
@@ 5 Conclusions
In this paper, we present DenseVLM, a framework designed to mitigate foreground bias in region-level vision-language alignment. DenseVLM can directly replace the original VLMs in open-vocabulary object detection and image segmentation methods, demonstrating consistently clear performance improvements over the baseline models. Furthermore, we validate DenseVLM’s promising scaling ability by exploring efficient region retrieval and decoupled alignment, successfully implementing DenseVLM to the training data from the SA-1B dataset. Overall, DenseVLM offers a generalizable solution for improving dense representations in vision-language models across various backbones, which is essential for open-vocabulary dense prediction tasks.
Acknowledgments
This work was partially funded by NSFC (No. 62225604, 62276145) and the Science and Technology Support Program of Tianjin, China (No. 23JCZDJC01050). This work was also supported by the Supercomputing Center of Nankai University.